# Functional Percolation: A Perspective on Criticality of Form and Function

### v2

@@ Abstract
Understanding the physical constraints and minimal conditions that enable information processing in extended systems remains a central challenge across disciplines, from neuroscience and artificial intelligence to social and physical networks. Here we study how network connectivity both limits and enables information processing by analyzing random networks across the structural percolation transition. Using cascade-mediated dynamics as a minimal and universal mechanism for propagating state-dependent responses, we examine structural, functional, and information-theoretic observables as functions of mean degree in Erdős–Rényi networks. We find that the emergence of a giant connected component coincides with a sharp transition in realizable information processing: complex input–output response functions become accessible, functional diversity increases rapidly, output entropy rises, and directed information flow—quantified by transfer entropy—extends beyond local neighborhoods. These coincident transitions define a regime of functional percolation, referring to a sharp expansion of the space of realizable input–output functions at the structural percolation transition. Near criticality, networks exhibit a Pareto-optimal tradeoff between functional complexity and diversity, suggesting that percolation criticality provides a universal organizing principle for information processing in systems with local interactions and propagating influences.
I Introduction
@@
It has been widely recognized that network cascades exhibit hallmarks of universal spreading processes, closely tied to percolation [bak1988self, sethna2001crackling, watts2002simple]. The universality of network cascades appears to arise from the fundamental nature of their minimal physical ingredients: State encoding by system elements, thresholding, information propagation, and energy input to drive the cascade process. These qualities can be found in many neuronal, social, biological, and physical systems. One may also reasonably conjecture that cascades form a universal physical substrate for information processing, as node response patterns to inputs constitute Boolean logic functions [Wilkerson2022]. Arguably, due to this dependence on very simple physical constraints and mechanisms, cascades may be more fundamental than other widely-studied Boolean network systems [Kauffman1969].
@@
Together, these points indicate that cascades and their induced Boolean transformations are generic computational responses in a large class of systems with local state, thresholds, and propagating interactions. They therefore provide a natural baseline for understanding how complex information processing emerges in physical and biological systems. Although we have not formalized an axiomatic framework here, the minimality and ubiquity of these ingredients suggest that cascades are promising candidates for developing a principled axiomatization of information processing in physical systems.
In this work, we make this intuition explicit. We show that while percolation in random networks enables large-scale information integration through the emergence of the giant connected component, the resulting Boolean response space grows combinatorially faster than the underlying configuration space of the system. This scale mismatch poses a fundamental challenge for information processing in physical networks, demanding that systems exploit all available dynamical and structural mechanisms. We demonstrate that criticality at the percolation transition provides a natural resolution of this challenge, simultaneously maximizing functional complexity, diversity, and information flow, a phenomenon we term functional percolation.
II The Core Problem: Scale Mismatch Between Form and Function
@@
III Criticality and Functional Percolation
In the universal information processing by cascades in random networks, outlined above, the ability of such systems to perform integrative information processing depends on large-scale connectivity [Wilkerson2022]. Integrative responses are not possible in fragmented random systems. The emergence of global connectivity comes in the form of structural percolation, with the emergence of the giant connected component (GCC) at the critical connectivity $z_{c}$ [Erdos1960RandomGraphs, Stauffer1994Percolation]. Thus the GCC size versus connectivity is a macroscopic order parameter.
Via cascades on the underlying emergent GCC, this symmetry breaking of form maps to a symmetry-breaking of function. The network topology, by realizing computational graphs, constrains the realizable functional responses. That is, there is a difference between the double-exponential abstract function space on $k$ inputs, and the inverse-exponentially scaling physically accessible function space. In this sense, functional percolation represents not merely the emergence of connectivity, but a phase transition in the realizable information processing of physical systems, arising due to the coincidence of structural and functional transitions.
@@ III Criticality and Functional Percolation
We studied linear-threshold cascades on Erdős–Rényi networks with $N=10^{4}$, homogeneous threshold $\phi=0.1$, and $k=5$ externally clamped input nodes, averaging functional and information-theoretic observables over independent realizations as a function of mean degree $z$. Across simulations, multiple measures of structure, function, and information processing exhibit pronounced changes at the structural percolation threshold $z_{c}\approx 1$ [Fig. 1].
@@
At criticality, the emergence of the giant connected component is accompanied by a scale-invariant distribution of finite components, which have been shown previously to act as small functional motifs in cascade dynamics [Wilkerson2022]. Due to the simultaneous growth of both DTC and functional diversity near criticality, we observe that this point represents a compromise at Pareto optimality between complexity and diversity of information processing [Langton1990EdgeOfChaos] [Fig. 3]. Exceeding the critical point suppresses exploratory capacity through over-constrained collective dynamics, while connectivity below criticality leads to fragmentation and an inability to realize complex functions. Thus, in the context of the overwhelming size of the Boolean function space, a system optimizes its exploration of diverse, complex functions by operating near criticality.
@@
These considerations motivate the interpretation of criticality as a pivot point between two distinct functional regimes: one supporting the execution of complex, integrative computations, and another enabling efficient exploration of function space. From this perspective, criticality may be viewed as a natural operating point for balancing exploration and exploitation, a distinction that may be framed phenomenologically in terms of reinforcement learning in adaptive or neural systems.
IV Discussion and Conclusion
@@
Our focus here has been on the physical precursors of information processing, framing them in terms of the most basic and universal physical processes. While these may and seem to underpin eventual Turing complete computation, or statistical models of learning, meriting further exploration in those contexts, we do not yet make those claims here. Rather, we begin from first principles of universal characteristics in physical systems to understand what emerges. We observe a universality of cascades in random networks as an information processing substrate, representing generic dynamics in thresholded, information propagating systems, underlying and independent of specifics found in biological, artificial, or social systems.
@@
From our results [Figure 1], we observe an agreement between functional and information-theoretic criticality. Transfer entropy can be viewed as a signature of information-flow percolation [schreiber2000measuring]. This gives support to criticality of information processing beyond Boolean functionality. This offers a structural-functional mechanism underlying the neural criticality hypothesis, that optimal information processing happens near criticality, giving us guidance of this behavior from first principles [BeggsPlenz2003, Shew2009DynamicRange, Kinouchi2006DynamicRange]. In the context of the scale mismatch between configuration space and function space discussed above, functional percolation thus appears as a natural way to reconcile limited structural resources with rich information-processing capabilities.
@@
Here we have observed how structural constraints can shape information states. Also, our results for functional diversity [Figure 1] suggest that, near criticality, the system can access a wide range of functions without needing extensive physical rewiring or many distinct configurations. Thus, remaining near criticality minimizes the structural ’work’ needed to explore information-processing states, because function-space exploration is maximized per unit structural change, and fewer physical modifications are needed to reach many computational states. Due to the compelling universal nature of the scale mismatch outlined above, this suggests that networks must use every possible tool at their disposal to efficiently explore the function space.
@@
We note that interaction topology and information propagation in networks allow us to begin to connect Boltzmann and Shannon entropies in a direct way, grounded in explicit causal interactions rather than ensemble-level statistical assumptions. At a deeper level, the simultaneous increase of functional complexity, entropy, and information flow at criticality reflects the removal of structural constraints at the percolation transition, which dramatically expands the accessible state space of the system; while this invites a thermodynamic interpretation in terms of effective entropy under coarse-graining, we do not attempt a formal entropy-balance or energetic treatment here. While we are not here making exact formal statements about Landauer bit-erasure or thermodynamics [Landauer1961Irreversibility], it is very compelling to undertake more formal exploration.
@@
Due to the compelling universal nature of the scale mismatch outlined above, one may expect that analogous optimizing mechanisms arise in other physical and biological systems.
@@
Finally, due to the basic and system-independent nature of the dynamics outlined here, we propose that functional percolation is a unifying physical principle. There is a coincidence of structural, functional, and informational phase transitions. While percolation theory is formally defined in the $N\to\infty$ limit, the sharp transitions observed here at large system size across multiple functional and information-theoretic order parameters are consistent with the existence of a well-defined functional percolation transition in the thermodynamic limit. In this sense, functional percolation is not only a useful operating regime, but a generic consequence of the structural and combinatorial constraints imposed on information processing in extended physical systems. This motivates extensive further investigation and formalization of the topics discussed here.

### v3

# Functional Percolation: Criticality of Form and Function

@@ Abstract
Understanding how network structure constrains and enables information processing is a central problem in statistical mechanics of interacting systems [Broadbent1957, Stauffer1994Percolation]. Here we study random networks across the structural percolation transition and analyze how connectivity governs realizable input–output transformations under cascade dynamics. Using Erdős–Rényi networks as a minimal ensemble, we examine structural, functional, and information-theoretic observables as functions of mean degree. We find that the emergence of the giant connected component coincides with a sharp transition in realizable information processing: complex input–output response functions become accessible, functional diversity increases rapidly, output entropy rises, and directed information flow—quantified by transfer entropy—extends beyond local neighborhoods. We term this coincidence of structural, functional, and informational transitions functional percolation, referring to a sharp expansion of the space of realizable input–output functions at the percolation threshold. Near criticality, networks exhibit a Pareto-optimal tradeoff between functional complexity and diversity, suggesting that percolation criticality may provide a general organizing principle of information processing capacity in systems with local interactions and propagating influences.
1 Introduction
@@
It has been widely recognized that network cascades exhibit hallmarks of universal spreading processes, closely tied to percolation [bak1988self, sethna2001crackling, watts2002simple]. The universality of network cascades appears to arise from the fundamental nature of their minimal physical ingredients: State encoding by system elements, thresholding, information propagation, and energy input to drive the cascade process. These qualities can be found in many neuronal, social, biological, and physical systems. One may also reasonably conjecture that cascades form a general physical substrate for information processing, as node response patterns to inputs constitute Boolean logic functions [Wilkerson2022]. Arguably, due to this dependence on very simple physical constraints and mechanisms, cascades may represent a particularly minimal and broadly applicable class of Boolean network dynamics [Kauffman1969, watts2002simple]. While microscopic interactions are continuous, coarse-graining generically produces state-dependent, threshold-like transitions, making cascade dynamics a widely applicable effective description of information propagation in interacting systems.
@@
Together, these points suggest that cascades and their induced Boolean transformations are generic computational responses in a large class of systems with local state, thresholds, and propagating interactions. They therefore provide a natural baseline for understanding how complex information processing emerges in physical and biological systems. Although we have not formalized an axiomatic framework here, the minimality and ubiquity of these ingredients suggest that cascades are promising candidates for developing a possible principled axiomatization of information processing in physical systems.
In this work, we make this intuition explicit. We show that while percolation in random networks enables large-scale information integration through the emergence of the giant connected component, the resulting Boolean response space grows combinatorially faster than the underlying configuration space of the system. This scale mismatch poses a fundamental challenge for information processing in physical networks, demanding that systems exploit all available dynamical and structural mechanisms. We demonstrate that criticality at the percolation transition provides a natural resolution of this challenge, simultaneously maximizing functional complexity, diversity, and information flow within this cascade model, a phenomenon we term functional percolation.
To our knowledge, this work provides systematic evidence that structural percolation induces a sharp transition in the space of realizable Boolean functions and associated information-theoretic observables.
2 The Core Problem: Scale Mismatch Between Form and Function
@@
3 Criticality and Functional Percolation
In the universal information processing by cascades in random networks, outlined above, the ability of such systems to perform integrative information processing depends on large-scale connectivity [Wilkerson2022]. Integrative responses are not possible in fragmented random systems. The emergence of global connectivity comes in the form of structural percolation, with the emergence of the giant connected component (GCC) at the critical connectivity $z_{c}$ [Erdos1960RandomGraphs, Stauffer1994Percolation]. Thus the GCC size versus connectivity is a macroscopic order parameter [Stanley1971].
Via cascades on the underlying emergent GCC, this symmetry breaking of form maps to a symmetry-breaking of function. The network topology, by realizing computational graphs, constrains the realizable functional responses. That is, there is a difference between the double-exponential abstract function space on $k$ inputs, and the inverse-exponentially scaling physically accessible function space. In this sense, functional percolation represents not merely the emergence of connectivity, but a phase transition-like change in the realizable information processing of physical systems, arising due to the coincidence of structural and functional transitions.
@@ 3.1 Model and observables
We studied linear-threshold cascades on Erdős–Rényi networks with $N=10^{4}$, homogeneous threshold $\phi=0.1$, and $k=5$ externally clamped input nodes, averaging functional and information-theoretic observables over independent realizations as a function of mean degree $z$. Across simulations, multiple measures of structure, function, and information processing exhibit pronounced changes at the structural percolation threshold $z_{c}\approx 1$ for fixed $k$ and fixed threshold parameters [Fig. 1].
@@
At criticality, the emergence of the giant connected component is accompanied by a scale-invariant distribution of finite components, which have been shown previously to act as small functional motifs in cascade dynamics [Wilkerson2022]. Due to the simultaneous growth of both DTC and functional diversity near criticality, we observe that this point appears to represent a compromise at Pareto optimality between complexity and diversity of information processing [Langton1990EdgeOfChaos] [Fig. 3]. Exceeding the critical point suppresses exploratory capacity through over-constrained collective dynamics, while connectivity below criticality leads to fragmentation and an inability to realize complex functions. Thus, in the context of the overwhelming size of the Boolean function space, a system optimizes its exploration of diverse, complex functions by operating near criticality.
@@
These considerations motivate the interpretation of criticality as a pivot point between two distinct functional regimes: one supporting the execution of complex, integrative computations, and another enabling efficient exploration of function space. From this perspective, criticality may be viewed as a natural operating point for balancing exploration and exploitation, a distinction that may be framed phenomenologically in terms of reinforcement learning in adaptive or neural systems [Fisher1967, Mora2011, KinouchiCopelli2006].
4 Discussion and Conclusion
@@
Our focus here has been on the physical precursors of information processing, framing them in terms of the most basic and universal physical processes. While these may and seem to underpin eventual Turing complete computation, or statistical models of learning, meriting further exploration in those contexts, we do not yet make those claims here. Rather, we begin from first principles of universal characteristics in physical systems to understand what emerges. We observe that cascades in random networks provide a generic information processing substrate, representing dynamics in thresholded, information propagating systems, underlying and independent of specifics found in biological, artificial, or social systems.
@@
From our results [Figure 1], we observe an agreement between functional and information-theoretic criticality. Transfer entropy can be provides a complementary indicator of directed dynamical influence [schreiber2000measuring]. This gives support to criticality of information processing beyond Boolean functionality. This offers a structural-functional mechanism underlying the neural criticality hypothesis, that optimal information processing happens near criticality, giving us guidance of this behavior from first principles [BeggsPlenz2003, Shew2009DynamicRange, Kinouchi2006DynamicRange]. In the context of the scale mismatch between configuration space and function space discussed above, functional percolation thus appears as a natural way to reconcile limited structural resources with rich information-processing capabilities.
@@
Here we have observed how structural constraints can shape information states. Also, our results for functional diversity [Figure 1] suggest that, near criticality, the system can access a wide range of functions without needing extensive physical rewiring or many distinct configurations. Thus, remaining near criticality minimizes the structural ’work’ needed to explore information-processing states, because function-space exploration is maximized per unit structural change, and fewer physical modifications are needed to reach many computational states. Due to the universal nature of the scale mismatch outlined above, this suggests that networks must use every possible tool at their disposal to efficiently explore the function space.
@@
We note that interaction topology and information propagation in networks allow us to begin to connect Boltzmann and Shannon entropies in a direct way, grounded in explicit causal interactions rather than ensemble-level statistical assumptions [sethna2001crackling, Jaynes1957]. At a deeper level, the simultaneous increase of functional complexity, entropy, and information flow at criticality reflects the removal of structural constraints at the percolation transition, which dramatically expands the accessible state space of the system; while this invites a thermodynamic interpretation in terms of effective entropy under coarse-graining, we do not attempt a formal entropy-balance or energetic treatment here. While we are not here making exact formal statements about Landauer bit-erasure or thermodynamics [Landauer1961Irreversibility], it is very compelling to undertake more formal exploration.
@@
Due to the compelling fundamental nature of the scale mismatch outlined above, one may expect that analogous optimizing mechanisms arise in other physical and biological systems.
@@
Finally, due to the basic and system-independent nature of the dynamics outlined here, we propose that functional percolation is a candidate unifying physical mechanism. There is a coincidence of structural, functional, and informational phase transitions. While percolation theory is formally defined in the $N\to\infty$ limit, the sharp transitions observed here at large system size across multiple functional and information-theoretic order parameters are consistent with the existence of a well-defined functional percolation transition in the thermodynamic limit. In this sense, functional percolation is not only a useful operating regime, but appears to be a generic consequence of the structural and combinatorial constraints imposed on information processing in extended physical systems. This motivates extensive further investigation and formalization of the topics discussed here.